OpenStudy (anonymous):

What is the exact value of √75/21? @jim_thompson5910 Can you help me out again please? lol.

jimthompson5910 (jim_thompson5910):

did you mean "exact value"?

OpenStudy (anonymous):

Yeah lol sorry, just finished a Geometry exam xD

jimthompson5910 (jim_thompson5910):

Can you factor 75 at all (where one factor is a perfect square)?

OpenStudy (anonymous):

Well 75 isn't a perfect square.

jimthompson5910 (jim_thompson5910):

that's true, but can you factor 75 into something like k*m where either k or m are perfect squares?

OpenStudy (anonymous):

23*3

jimthompson5910 (jim_thompson5910):

you mean 25*3?

OpenStudy (anonymous):

yeah. man I can't type today lol but yes that's what I meant.

jimthompson5910 (jim_thompson5910):

so this means... sqrt(75) = sqrt(25*3) sqrt(75) = sqrt(25)*sqrt(3) sqrt(75) = 5*sqrt(3)

jimthompson5910 (jim_thompson5910):

So \[\Large \frac{\sqrt{75}}{21} = \frac{5\sqrt{3}}{21}\]

OpenStudy (anonymous):

okay, so now what happens with the 21?

jimthompson5910 (jim_thompson5910):

wait a second, is the 21 in the square root as well?

OpenStudy (anonymous):

yes it is

jimthompson5910 (jim_thompson5910):

alright, that completely changes things

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

thanks

OpenStudy (anonymous):

Sorry I should have made that a bit clearer in the beginning.

jimthompson5910 (jim_thompson5910):

that's fine

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{\frac{75}{21}}\] \[\Large \sqrt{\frac{25*3}{7*3}}\] \[\Large \sqrt{\frac{25}{7}}\] \[\Large \frac{\sqrt{25}}{\sqrt{7}}\] \[\Large \frac{5}{\sqrt{7}}\] \[\Large \frac{5*\sqrt{7}}{\sqrt{7}*\sqrt{7}}\] \[\Large \frac{5*\sqrt{7}}{\sqrt{7*7}}\] \[\Large \frac{5*\sqrt{7}}{\sqrt{49}}\] \[\Large \frac{5\sqrt{7}}{7}\] So \[\Large \sqrt{\frac{75}{21}} = \frac{5\sqrt{7}}{7}\]

OpenStudy (anonymous):

Oh okay. So where did you get the 7 * 3 from?

jimthompson5910 (jim_thompson5910):

21 = 7*3

jimthompson5910 (jim_thompson5910):

I factored each with a factor of 3 so the 3's will cancel

OpenStudy (anonymous):

ahhh okay that makes sense. Thank you again!!